OpenStudy (goldrush18):

differentiate with respect to x: i. y=cos(2x+5)+2xsin3x ii. y=3x^2-1/x^2+1

OpenStudy (lalaly):

have you tried to do it?

OpenStudy (goldrush18):

yes but i'm not sure its correct

OpenStudy (lalaly):

ok can u post your answers coz if theyr wrong ill show you how to do it

OpenStudy (goldrush18):

i. y^1= -2 sin 2x+5+6x cos 3x

OpenStudy (anonymous):

ooh i see lalaly is already on it :P

OpenStudy (anonymous):

she's pro

OpenStudy (lalaly):

well thats right but u forgot the product rule in the next part y'= -2sin(2x+5)+6xcos(3x)+2sin(3x) because 2xsin(3x) is two functions multiplied by eachother so when u want to differentiate that u use the product rule

OpenStudy (lalaly):

lol tai thanks xD

OpenStudy (goldrush18):

im working on the next one

OpenStudy (lalaly):

oki im waiting for ur answer:)

OpenStudy (anonymous):

\[dy/dx=6x+2x^-3\]

OpenStudy (anonymous):

2x^-3?

OpenStudy (anonymous):

wait no! lol

OpenStudy (goldrush18):

\[y=3x^2-1\div x^2+1\] \[=6x(x^2+1)-2x(3x^2-1)\div(x^2+1)^2\] \[=6x^3+6x-6x^3+2x \div (x^2+1)^2\] \[y^1=8x \div (x^2+1)^2\]

OpenStudy (anonymous):

goldrush, put brackets please

OpenStudy (lalaly):

yup u got it ;) @goldrush18

OpenStudy (anonymous):

yeah that's it

OpenStudy (goldrush18):

omg really

OpenStudy (anonymous):

wait checking again

OpenStudy (lalaly):

yes its perfect

OpenStudy (goldrush18):

awesome well i have other questions coming up that i dont understand they are on LIMITS

OpenStudy (anonymous):

you should state when you perform dy/dx to show examiner that you know

OpenStudy (goldrush18):

dy/dx or y^1 thats how we do it as well

OpenStudy (anonymous):

it's prime :P not ^1 haha

OpenStudy (lalaly):

\[y'\]\[y^{(1)}\]and\[\frac{dy}{dx}\] are all the same tai

OpenStudy (goldrush18):

i have exams coming up soon so i have to get the hang of these questions

OpenStudy (anonymous):

wait it's ^1?

OpenStudy (anonymous):

is this your last year btw goldrush?

OpenStudy (lalaly):

gold rush im here if u need any help just tag me:)

OpenStudy (goldrush18):

@Tai yes it is

OpenStudy (anonymous):

nice, i have one more year after this

OpenStudy (anonymous):

I have soo many exams... lol

OpenStudy (goldrush18):

thanks @lalaly i do both u and @Tai

OpenStudy (anonymous):

SAT1, SSAT Maths2, SSAT physics, SSAT chemistry, AS Maths C1,C2,S1, Macroeconomics, Microeconomics, Physics U1, U2, Practical, Chemistry U1,U2, Practical

OpenStudy (anonymous):

15 exams lol

OpenStudy (goldrush18):

wow thats alot to study for!!

OpenStudy (goldrush18):

well i have Pure Mathematics, Accounting, Caribbean Studies and Management of Business

OpenStudy (anonymous):

:O Caribbean studies? what's that?

OpenStudy (goldrush18):

for now until i move on to University

OpenStudy (anonymous):

are you studying to become a pirate? :P

OpenStudy (goldrush18):

its basically the whole history about the Caribbean

OpenStudy (goldrush18):

No lol! its a compulsory subject

OpenStudy (anonymous):

and how to become a pirate :]